# Neutrino flavor instabilities in neutron star mergers with moment transport: slow, fast, and collisional modes

### v1

@@ Abstract
Determining where, when, and how neutrino flavor oscillations must be included in large-scale simulations of hot and dense astrophysical environments is an enduring challenge that must be tackled to obtain accurate predictions. Using an angular moment-based linear stability analysis framework, we examine the different kinds of flavor instabilities that can take place in the context of the post-processing of a neutron star merger simulation, with a particular focus on the collisional flavor instability and a careful assessment of several commonly used approximations. First, neglecting anisotropies of the neutrino field, we investigate the extent to which commonly used monoenergetic growth rates reproduce the results obtained from a full multi-energy treatment. Contrary to the large discrepancies found in core-collapse supernova environments, we propose a simple combination of energy-averaged estimates that reproduces the multi-energy growth rates in our representative simulation snapshot. We then quantify the impact of additional physical effects, including nuclear many-body corrections, scattering opacities, and the inclusion of the vacuum term in the neutrino Hamiltonian. Finally, we include the neutrino distribution anisotropies, which allows us to explore, for the first time in a multi-energy setting, the interplay between collisional, fast, and slow modes in a moment-based neutron star merger simulation. We find that despite a dominance of the fast instability in most of the simulation volume, certain regions only exhibit a collisional instability, while others, especially at large distances, exhibit a slow instability that is largely underestimated if anisotropic effects are neglected.
@@ I Introduction
To determine whether neutrino flavor transformation can occur for a given flavor configuration, it is natural to perform a linear stability analysis (LSA) [48]. This analysis requires linearizing the quantum kinetic equations (QKEs), ideally written for all of the system’s momentum degrees of freedom: the different propagation directions (angular distribution) and the magnitudes (energies). We base our LSA on the framework of Ref. [29], which derived the LSA equations in the form of a stability matrix eigenvalue problem for a single-energy system, and employed only the first few angular moments in place of the full angular distribution. We generalize this formalism by considering a multi-energy spectrum, and include the vacuum and collision terms from the QKEs. We frequently discuss the different unstable modes and their characteristics in light of the recent comprehensive study by Fiorillo and Raffelt [49], which uses an equivalent formalism to the stability matrix method, the so-called “dispersion relation approach” [50].
@@
Some recent studies [39, 29, 51] have looked at the occurrence of CFI in classical postmerger simulation snapshots, and they all estimated the growth rate assuming that the neutrino distributions were isotropic, an assumption that fails beyond the neutrino decoupling region. Furthermore, it is common in these studies to invoke growth rate formulas for a monochromatic gas using energy-averaged collision rates, following analytic and numerical arguments in [38, 41]. It was, however, shown recently in [47] that using energy-averaged collision rates dramatically overestimates the occurrence of CFIs in CCSNe. In this paper we examine this same, energy-averaged, method of stability analysis using a NSM simulation snapshot from [52], and find that the different properties of the flavor-dependent neutrino distributions in CCSNe and NSMs actually allow for a good energy-averaged description in the latter. Our work aims at providing a complete picture of the possible instabilities that can occur in a NSM postmerger environment, describing slow, collisional and fast modes.
@@
A LSA informs us of the locations and local properties (wavelength, timescale) of flavor instabilities, but not on their eventual, asymptotic, outcome. This aspect is beyond the scope of this paper, but is being intensely investigated for fast instabilities [53, 54, 55, 56, 57, 58, 59, 60, 61, 62], which has allowed for the integration of subgrid models of the FFI in large-scale simulations of CCSNe [63, 64] and postmerger disks [65]. The outcome of the CFI has been less intensely studied, see e.g., [66, 67].
@@
Our paper is organized as follows. In Sec. II, we introduce the neutrino evolution equations and their version written for angular moments, which we linearize to derive our LSA framework. For the specific case of CFIs with isotropic neutrino distributions, we discuss various monochromatic descriptions of the system (see also Appendix A for useful analytic formulas). We discuss in Appendix B the connection of our LSA method, based on a stability matrix, with the commonly-used “dispersion relation” approach. We introduce in Sec. III the NSM simulation data we study in this work, outlining the approximations we make to obtain the angular and energy neutrino distributions at each simulation grid point. We then describe the neutrino flavor instabilities occurring in this NSM snapshot, first assuming in Sec. IV, for the sake of comparison with the CFI literature, that neutrino distributions are isotropic. We notably discuss how energy-averaged methods can provide a good estimate of the CFI growth rate in this environment, with a study of the difference with CCSN configurations in Appendix C. We investigate the changes in the CFI landscape when including nucleon many-body corrections to the absorption rates (Sec. IV.2), scattering opacities (Sec. IV.3), or the vacuum term (Sec. IV.4). In Sec. V, we study the flavor instabilities in the NSM snapshot when considering the actual neutrino anisotropies, which allows us to describe fast, slow and collisional instabilities. Our findings are confirmed in other NSM snapshots, which we discuss in Appendix D. Finally, we summarize and conclude in Sec. VI.
@@ II.1 Quantum Kinetic Equations
The statistical ensemble of (anti)neutrinos, including flavor mixing, are described by one-body density matrices for neutrinos $\varrho(t,\mathbf{x},\mathbf{p})$] and antineutrinos $\bar{\varrho}(t,\mathbf{x},\mathbf{p})$], where $t$ is time, $\mathbf{x}$ is the spatial coordinates, and $\mathbf{p}$ is the (anti)neutrino momentum. They are matrices in flavor space, where the on-diagonal components generalize the classical distribution functions (we will sometimes use the notation $f_{\nu_{\alpha}}=\varrho_{\alpha\alpha}$), and the complex off-diagonal components measure the so-called “flavor coherence.” Their evolution is given by the QKEs [68, 69, 70, 71]
@@ II.2 Moment formalism
which we will respectively call “number density,” “(number) flux density,” and “pressure tensor.” It should be noted that $P_{\alpha\beta}^{ij}$ does not have here the units of a pressure, but we still use this name for convenience. The energy-integrated moments are
@@
In order to obtain a genuine two-moment scheme, that is, evolving only $(N,\overline{N})$ and $(\mathbf{F},\mathbf{\overline{F}})$, one needs to express the pressure tensors as a function of the first two moments. In particular, since flavor evolution can lead to a transfer of power to smaller angular scales [82, 83], this emphasizes the need for a proper closure. Although the closure problem for classical neutrino transport has been intensely studied (see, e.g., [84, 85, 86, 87, 88, 89]), the generalization to quantum moments has only been recently explored, see [90, 91]. In this work, we perform a linear stability analysis restricted to a specific wavenumber, for which the closure relation for the flavor off-diagonal components of the moments plays no role, and we leave a dedicated study of the impact and adequacy of the chosen quantum closure relation for future work.
@@ II.4 Linear stability analysis
We emphasize that the specific choice $\mathbf{k}^{\prime}=\mathbf{0}$ reduces the stability matrix dependence on the pressure tensors to their flavor on-diagonal components, which appear in $\Delta_{P}^{ij}$, $\overline{\Delta}_{P}^{ij}$. For other values of $\mathbf{k}^{\prime}$, one needs to also specify the closure for the off-diagonal component $P_{ex}^{ij}$, see Ref. [29]. The NSM simulation we study in this work uses the classical M1 two-moment scheme, such that we can borrow its closure prescription for $P_{\alpha\alpha}^{ij}$ [see Eq. (25)].
@@ II.5 Energy-averaged approaches
Much of the work on CFIs has focused on monochromatic energy distributions, or at least used the results of the monochromatic dispersion relation with certain energy-averaged quantities (e.g., [35, 37, 43, 44, 77, 67]). It is an attractive approach as it avoids the numerical complexity of the multi-energy analysis: for instance, it reduces by a factor $n_{\mathrm{groups}}^{2}$ the size of the stability matrix [Eq. (14)]. Such methods were justified by analytical and numerical results [38, 39, 41], which showed that the instability growth rate in multi-energy analyses was well recovered by monochromatic formulae.
@@ II.5.1 Gapped modes and energy-averaging “method A”
One can therefore estimate the multi-energy isotropic CFI growth rate by using the full monochromatic formulae (43) with the averaged collision rates (19). This energy-averaging approach is called “Method A” in [47], and we adopt this nomenclature. Note that, as demonstrated above, the gapped modes should be correctly described by this method, but nothing guarantees that the gapless modes would be well recovered. Indeed, Wang et al. have shown that the growth rates of gapless modes are significantly overestimated, even predicting instability when the system is actually stable. We return to this point in Sec. IV.
@@ II.5.2 Energy-averaging “method B”
Another monochromatic description, used for instance in [41, 44], consists of averaging the flavor-dependent opacities $\kappa_{\alpha}$ (we omit the “absorption” subscript), and then reconstructing the effective flavor coherence damping rate. Specifically, we define
@@ III.1 NSM simulation snapshots
We apply our linear stability analysis to the results from a general relativistic simulation of the merger of two neutron stars ($1.3\,M_{\odot}$ and $1.4\,M_{\odot}$) which uses a gray two-moment scheme for neutrino transport [96, 92], specifically the “M1-NuLib” simulation in [52]. We examine two different stages of the postmerger evolution, selecting snapshots at $3\,\mathrm{ms}$ (“early”) and $7\,\mathrm{ms}$ (“late”) post-merger—in the simulation, the hypermassive neutron star (HMNS) remnant collapses into a black hole after $8.5\,\mathrm{ms}$. In the main text, we focus on the $7\,\mathrm{ms}$ snapshot, with results for the $3\,\mathrm{ms}$ snapshot shown in Appendix D (along with a snaphot from an independent simulation, taken from Ref. [92]). We show on Fig. 1 some data from the simulation, for a vertical slice passing through the “late” snapshot of the remnant. The top-right panel of Fig. 1 shows density contours for this $7\,\mathrm{ms}$ postmerger snapshot. At that time, the HMNS is surrounded by an accretion torus of density $10^{10-12}\,{\rm g\,cm^{-3}}$. Within the torus are shocked tidal arms where the fluid temperature is higher and large density jumps can be seen. The tidal arms create sharp features in many of the vertical slices shown here (e.g., the region of high temperature $T_{\nu_{x}}$ at the density contour $\rho=10^{11}\,{\rm g\,cm^{-3}}$ on Fig. 1). The polar regions above and below the remnant are filled with low density outflowing matter.
@@ III.2 Neutrino distributions
The NSM simulation from which we extract the thermodynamics and neutrino data uses an energy-integrated two-moment scheme, from which we obtain the energy flux, energy density and number density in the inertial frame. Transforming these radiation field quantities into a frame comoving with the fluid (see, e.g., Appendix B in [98]), we can get the quantities $\mathcal{N}_{\alpha\alpha}$ and $\mathcal{F}_{\alpha\alpha}^{i}$ [see definitions in Eq. (6)]. To perform our linear stability analysis, we need the energy distributions $N_{\alpha\alpha}(p)$ and $F_{\alpha\alpha}^{i}(p)$, and the flavor-diagonal pressure tensor $P_{\alpha\alpha}^{ij}$. In addition, this simulation considers three neutrino species: $\nu_{e}$, $\bar{\nu}_{e}$ and a combination of the heavy-lepton flavor neutrinos and antineutrinos, assumed to have the same distributions, $\nu_{X}$. In our two-flavor analysis, we thus take the values for $\nu_{e}$, $\bar{\nu}_{e}$ and we assign schematically the distribution of a single heavy-lepton neutrino species $f_{\nu_{x}}=f_{\nu_{X}}/4$, $f_{\bar{\nu}_{x}}=f_{\nu_{X}}/4$. On the left panels of Fig. 1, we show ratios of the neutrino number densities $\mathcal{N}_{\alpha\alpha}$, comparing $\nu_{e}$ and $\bar{\nu}_{e}$ in the top panel, and $\nu_{x}$ and $\nu_{e}$ in the bottom panel.
@@
We recall that we use a hat over a vector to indicate the associated unit vector. The angular part of $f_{\nu_{\alpha}}$ is given by the classical Maximum Entropy (ME), or Minerbo, closure [99, 84, 87],
@@
The energy spectrum is assumed to be a gray Fermi-Dirac distribution,
@@
where the parameters $\{g_{\nu_{\alpha}},T_{\nu_{\alpha}}\}$ are determined by the two equations
@@
On the top right panel of Fig. 1, we show the ratio of the temperatures $T_{\nu_{\alpha}}$ entering the gray spectrum (27) for $\alpha=e,x$. We also show some matter density contours (see caption and discussion at the beginning of Sec. III). Finally, in the bottom right panel is shown which of the neutrinos or antineutrinos have the largest collision rate (restricting to absorption processes), focusing on the flavor-averaged collision rates (21). In the large majority of the slice, neutrinos have larger interaction rates, such that the gapless or gapped nature of CFIs is determined by which of $\nu_{e}$ or $\bar{\nu}_{e}$ is the most abundant (see top left panel).

### v2

@@ Abstract
Determining where, when, and how neutrino flavor oscillations must be included in large-scale simulations of hot and dense astrophysical environments is an enduring challenge that must be tackled to obtain accurate predictions. Using an angular moment-based linear stability analysis framework, we examine the different kinds of flavor instabilities that can take place in the context of the post-processing of a neutron star merger simulation, with a particular focus on the collisional flavor instability and a careful assessment of several commonly used approximations. First, neglecting anisotropies of the neutrino field, we investigate the extent to which commonly used monoenergetic growth rates reproduce the results obtained from a full multi-energy treatment. Contrary to the large discrepancies found in core-collapse supernova environments, we propose a simple combination of energy-averaged estimates that reproduces the multi-energy growth rates in our representative simulation snapshot. We then quantify the impact of additional physical effects, including nuclear many-body corrections, scattering opacities, and the inclusion of the vacuum term in the neutrino Hamiltonian. Finally, we include the neutrino distribution anisotropies, which allows us to explore, for the first time in a multi-energy setting, the interplay between collisional, fast, and slow modes in a moment-based neutron star merger simulation. We find that, despite a dominance of the fast instability in most of the simulation volume, certain regions exhibit only a collisional instability, while others, especially at large distances, exhibit a slow instability that is largely underestimated if anisotropic effects are neglected.
@@ I Introduction
To determine whether neutrino flavor transformation can occur for a given flavor configuration, it is natural to perform a linear stability analysis (LSA) [48]. This analysis requires linearizing the quantum kinetic equations (QKEs), ideally written for all of the system’s momentum degrees of freedom: the different propagation directions (angular distribution) and the magnitudes (energies). We base our LSA on the framework of Ref. [29], which derived the LSA equations in the form of a stability matrix eigenvalue problem for a single-energy system and employed only the first few angular moments in place of the full angular distribution. We generalize this formalism by considering a multi-energy spectrum, and include the vacuum and collision terms from the QKEs. We frequently discuss the different unstable modes and their characteristics in light of the recent comprehensive study by Fiorillo and Raffelt [49], which uses an equivalent formalism to the stability matrix method, the so-called “dispersion relation approach” [50].
@@
Some recent studies [39, 29, 51] have looked at the occurrence of CFI in classical postmerger simulation snapshots, and they all estimated the growth rate assuming that the neutrino distributions were isotropic, an assumption that fails beyond the neutrino decoupling region. Furthermore, it is common in these studies to invoke growth rate formulas for a monochromatic gas using energy-averaged collision rates, following analytic and numerical arguments in [38, 41]. It was, however, shown recently in [47] that using energy-averaged collision rates dramatically overestimates the occurrence of CFIs in CCSNe. In this paper we examine this same, energy-averaged, method of stability analysis using a NSM simulation snapshot from [52], and find that the different properties of the flavor-dependent neutrino distributions in CCSNe and NSMs actually allow for a good energy-averaged description in the latter. Our work aims at providing a complete picture of the possible instabilities that can occur in a NSM postmerger environment, describing slow, collisional, and fast modes.
@@
A LSA informs us of the locations and local properties (wavelength, timescale) of flavor instabilities, but not on their eventual, asymptotic, outcome. This aspect is beyond the scope of this paper but is being intensely investigated for fast instabilities [53, 54, 55, 56, 57, 58, 59, 60, 61, 62], which has allowed for the integration of subgrid models of the FFI in large-scale simulations of CCSNe [63, 64, 65] and postmerger disks [66]. The outcome of the CFI has been less intensely studied; see, e.g., [67, 68].
@@
Our paper is organized as follows. In Sec. II, we introduce the neutrino evolution equations and their version written for angular moments, which we linearize to derive our LSA framework. For the specific case of CFIs with isotropic neutrino distributions, we discuss various monochromatic descriptions of the system (see also Appendix A for useful analytic formulas). We discuss in Appendix B the connection of our LSA method, based on a stability matrix, with the commonly used “dispersion relation” approach. We introduce in Sec. III the NSM simulation data we study in this work, outlining the approximations we make to obtain the angular and energy neutrino distributions at each simulation grid point. We then describe the neutrino flavor instabilities occurring in this NSM snapshot, first assuming in Sec. IV, for the sake of comparison with the CFI literature, that neutrino distributions are isotropic. We notably discuss how energy-averaged methods can provide a good estimate of the CFI growth rate in this environment, with a study of the difference with CCSN configurations in Appendix C. We investigate the changes in the CFI landscape when including nucleon many-body corrections to the absorption rates (Sec. IV.2), scattering opacities (Sec. IV.3), or the vacuum term (Sec. IV.4). In Sec. V, we study the flavor instabilities in the NSM snapshot when considering the actual neutrino anisotropies, which allows us to describe fast, slow, and collisional instabilities. Our findings are confirmed in other NSM snapshots, which we discuss in Appendix D. Finally, we summarize and conclude in Sec. VI.
@@ II.1 Quantum Kinetic Equations
The statistical ensemble of (anti)neutrinos, including flavor mixing, is described by one-body density matrices for neutrinos $\varrho(t,\mathbf{x},\mathbf{p})$] and antineutrinos $\bar{\varrho}(t,\mathbf{x},\mathbf{p})$], where $t$ is time, $\mathbf{x}$ is the spatial coordinates, and $\mathbf{p}$ is the (anti)neutrino momentum. They are matrices in flavor space, where the on-diagonal components generalize the classical distribution functions (we will sometimes use the notation $f_{\nu_{\alpha}}=\varrho_{\alpha\alpha}$), and the complex off-diagonal components measure the so-called “flavor coherence.” Their evolution is given by the QKEs [69, 70, 71, 72]
@@ II.2 Moment formalism
which we will, respectively, call “number density,” “(number) flux density,” and “pressure tensor.” It should be noted that $P_{\alpha\beta}^{ij}$ does not have here the units of a pressure, but we still use this name for convenience. The energy-integrated moments are
@@
In order to obtain a genuine two-moment scheme, that is, evolving only $(N,\overline{N})$ and $(\mathbf{F},\mathbf{\overline{F}})$, one needs to express the pressure tensors as a function of the first two moments. In particular, since flavor evolution can lead to a transfer of power to smaller angular scales [83, 84], this emphasizes the need for a proper closure. Although the closure problem for classical neutrino transport has been intensely studied (see, e.g., [85, 86, 87, 88, 89, 90]), the generalization to quantum moments has only been recently explored; see [91, 92]. In this work, we perform a linear stability analysis restricted to a specific wavenumber, for which the closure relation for the flavor off-diagonal components of the moments plays no role, and we leave a dedicated study of the impact and adequacy of the chosen quantum closure relation for future work.
@@ II.4 Linear stability analysis
We emphasize that the specific choice $\mathbf{k}^{\prime}=\mathbf{0}$ reduces the stability matrix dependence on the pressure tensors to their flavor on-diagonal components, which appear in $\Delta_{P}^{ij}$, $\overline{\Delta}_{P}^{ij}$. For other values of $\mathbf{k}^{\prime}$, one needs to also specify the closure for the off-diagonal component $P_{ex}^{ij}$; see Ref. [29]. The NSM simulation we study in this work uses the classical M1 two-moment scheme, such that we can borrow its closure prescription for $P_{\alpha\alpha}^{ij}$ [see Eq. (25)].
@@ II.5 Energy-averaged approaches
Much of the work on CFIs has focused on monochromatic energy distributions, or at least used the results of the monochromatic dispersion relation with certain energy-averaged quantities (e.g., [35, 37, 43, 44, 78, 68]). It is an attractive approach as it avoids the numerical complexity of the multi-energy analysis: for instance, it reduces by a factor $n_{\mathrm{groups}}^{2}$ the size of the stability matrix [Eq. (14)]. Such methods were justified by analytical and numerical results [38, 39, 41], which showed that the instability growth rate in multi-energy analyses was well recovered by monochromatic formulas.
@@ II.5.1 Gapped modes and energy-averaging “method A”
One can therefore estimate the multi-energy isotropic CFI growth rate by using the full monochromatic formulas (43) with the averaged collision rates (19). This energy-averaging approach is called “Method A” in [47], and we adopt this nomenclature. Note that, as demonstrated above, the gapped modes should be correctly described by this method, but nothing guarantees that the gapless modes would be well recovered. Indeed, Wang et al. have shown that the growth rates of gapless modes are significantly overestimated, even predicting instability when the system is actually stable. We return to this point in Sec. IV.
@@ II.5.2 Energy-averaging “method B”
Another monochromatic description, used for instance in [41, 44], consists of averaging the flavor-dependent opacities $\kappa_{\alpha}$ (we omit the “absorption” subscript) and then reconstructing the effective flavor coherence damping rate. Specifically, we define
@@ III.1 NSM simulation snapshots
We apply our linear stability analysis to the results from a general relativistic simulation of the merger of two neutron stars ($1.3\,M_{\odot}$ and $1.4\,M_{\odot}$) which uses a gray two-moment scheme for neutrino transport [97, 93], specifically the “M1-NuLib” simulation in [52]. We examine two different stages of the postmerger evolution, selecting snapshots at $3\,\mathrm{ms}$ (“early”) and $7\,\mathrm{ms}$ (“late”) post-merger—in the simulation, the hypermassive neutron star (HMNS) remnant collapses into a black hole after $8.5\,\mathrm{ms}$. In the main text, we focus on the $7\,\mathrm{ms}$ snapshot, with results for the $3\,\mathrm{ms}$ snapshot shown in Appendix D (along with a snapshot from an independent simulation, taken from Ref. [93]). We show in Fig. 1 some data from the simulation, for a vertical slice passing through the “late” snapshot of the remnant. The top-right panel of Fig. 1 shows density contours for this $7\,\mathrm{ms}$ postmerger snapshot. At that time, the HMNS is surrounded by an accretion torus of density $10^{10-12}\,{\rm g\,cm^{-3}}$. Within the torus are shocked tidal arms where the fluid temperature is higher and large density jumps can be seen. The tidal arms create sharp features in many of the vertical slices shown here (e.g., the region of high temperature $T_{\nu_{x}}$ at the density contour $\rho=10^{11}\,{\rm g\,cm^{-3}}$ on Fig. 1). The polar regions above and below the remnant are filled with low-density outflowing matter.
@@ III.2 Neutrino distributions
The NSM simulation from which we extract the thermodynamics and neutrino data uses an energy-integrated two-moment scheme, from which we obtain the energy flux, energy density and number density in the inertial frame. Transforming these radiation field quantities into a frame comoving with the fluid (see, e.g., Appendix B in [99]), we can get the quantities $\mathcal{N}_{\alpha\alpha}$ and $\mathcal{F}_{\alpha\alpha}^{i}$ [see definitions in Eq. (6)]. To perform our linear stability analysis, we need the energy distributions $N_{\alpha\alpha}(p)$ and $F_{\alpha\alpha}^{i}(p)$, and the flavor-diagonal pressure tensor $P_{\alpha\alpha}^{ij}$. In addition, this simulation considers three neutrino species: $\nu_{e}$, $\bar{\nu}_{e}$ and a combination of the heavy-lepton flavor neutrinos and antineutrinos, assumed to have the same distributions, $\nu_{X}$. In our two-flavor analysis, we thus take the values for $\nu_{e}$ and $\bar{\nu}_{e}$ and we assign schematically the distribution of a single heavy-lepton neutrino species $f_{\nu_{x}}=f_{\nu_{X}}/4$, $f_{\bar{\nu}_{x}}=f_{\nu_{X}}/4$. In the left panels of Fig. 1, we show ratios of the neutrino number densities $\mathcal{N}_{\alpha\alpha}$, comparing $\nu_{e}$ and $\bar{\nu}_{e}$ in the top panel and $\nu_{x}$ and $\nu_{e}$ in the bottom panel.
@@
We recall that we use a hat over a vector to indicate the associated unit vector. The angular part of $f_{\nu_{\alpha}}$ is given by the classical maximum entropy (ME), or Minerbo, closure [100, 85, 88]:
@@
The energy spectrum is assumed to be a gray Fermi-Dirac distribution:
@@
where the parameters $\{g_{\nu_{\alpha}},T_{\nu_{\alpha}}\}$ are determined by the two equations:
@@
In the top right panel of Fig. 1, we show the ratio of the temperatures $T_{\nu_{\alpha}}$ entering the gray spectrum (27) for $\alpha=e,x$. We also show some matter density contours (see caption and discussion at the beginning of Sec. III). Finally, in the bottom right panel is shown which of the neutrinos or antineutrinos have the largest collision rate (restricting to absorption processes), focusing on the flavor-averaged collision rates (21). In the large majority of the slice, neutrinos have larger interaction rates, such that the gapless or gapped nature of CFIs is determined by which of $\nu_{e}$ or $\bar{\nu}_{e}$ is the most abundant (see top left panel).